Grade 11General Physics

a particle of mass m1 is fastened to one end of a string and one of m2 to the middle point,the other end of the string being fastened to a fixed point on a smooth horizontal table. The particles are then projected,so that the two portions of the strings are always in the same straight line and describes horizontal circles.Find the ratio of tensions in the two parts of the string.

khushi
10 Years agoGrade 11

1 Answer

Saloni Doshi
8 Years ago
For particle with mass m1, 1/2m1(v1)2= 2m1g(2h) [ according to conservation of mechanical energy] therefore v1=(8gh)1/2.     Since the tension in the string provides the centripetal force therefore T1=m1(v1)2/2h   We get T= 4m1gh.   For particle with mass m2, 1/2m2(v2)2=   2m2gh therefore v2= (4gh)1/2.   For this particle T2-T1=m2(v2)2/2h = 2m2g therefore from these equations we get T2= 2m2g + 4m1g   Therefore T1/T2= 2m1/m2+2m1